### v1

# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Bridges

Abstract
Between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen via attacks on “bridges” — decentralized services designed to allow inter-blockchain exchange. While the individual exploits in each attack vary, a single design flaw underlies them all: the lack of end-to-end value accounting in cross-chain transactions. In this paper, we empirically analyze twenty million transactions used by key bridges during this period. We show that a simple invariant that balances cross-chain inflows and outflows is compatible with legitimate use, yet precisely identifies every known attack (and several likely attacks) in this data. Further, we show that this approach is not only sufficient for post-hoc audits, but can be implemented in-line in existing bridge designs to provide generic protection against a broad array of bridge vulnerabilities.
1 Introduction
Careful accounting is key to the correct management of assets in virtually all financial systems. Indeed, since Paciloli popularized double-entry bookkeeping in the 14th century, it has been standard practice to separately track inflows (credits) and outflows (debits) to establish a net position — a balance sheet.
Such accounting is implicit in blockchains as each and every transaction is recorded, immutable and in possession of implicit integrity. Any change in ownership for a given token is explicitly recorded via some past transaction and thus the net position for assets in a blockchain is generally consistent and well-known. However, these integral properties only hold within a single blockchain. As soon as one wishes to engage in transactions between blockchains (e.g., trading Ethereum on Solana), it requires building a system that steps outside these isolated environments and implements its own financial calculus between them, including its own accounting.
Today, one of the principal mechanisms for such transactions is the cross-chain bridge — a service using a pair of “smart contracts” (immutable programs stored on a blockchain) to synthesize inter-chain transactions that are not possible to express natively on a single chain. [^1] Some newer blockchain ecosystems, such as Cosmos and Avalanche, are themselves composed of multiple blockchains and offer some support for bridging across their own chains, but not with external chains. Such bridges are an extremely popular component of what is commonly referred to as “decentralized finance” (DeFi), and they routinely manage transaction volumes with value equivalent to \$US8–10 billion per month [7].
However, bridges are just code and they can have bugs in their implementations, in the other services they depend upon, or in the mechanisms used to guard the integrity of their cryptographic secrets. As a result, attackers can, and do, exploit such vulnerabilities to extract significant value from such systems. For example, between 2021 and 2023, crypto assets valued at over \$US2.6 billion were stolen in an array of attacks on bridges. In this work, we identify the lack of explicit accounting as the shared and signal failure across virtually all of these attacks. Considering the twelve largest attacks (each \$US1m or more) during this period, the details of their vulnerabilities vary considerably, but all share the property that transactions are allowed to be “unbalanced” (i.e., that outflow can exceed inflow, less costs). We show that auditing transactions for this property is sufficient to retrospectively and automatically identify each of these past attacks. Moreover, such audits can be performed in real-time — either to alert bridge operators about potential fraud, or prevent such unbalanced transactions from completing altogether.
In particular, our paper makes three main contributions:
  • $\triangleright$
    Balance Invariant. We introduce the notion of a simple balance invariant, outflow = inflow - costs, which can be calculated automatically from the transaction formats used by current bridges. We hypothesize that this invariant should naturally hold for benign transactions, but not for those exploiting bridge vulnerabilities.
  • $\triangleright$
    Retrospective validation. We validate our hypothesis by empirically auditing over 20 million past transactions (across more than 20 blockchains, between 11 different bridges) and show that the balance invariant is sufficient to identify each and every attack for which we have collected ground truth. Moreover, we identify vanishingly few other transactions that violate this invariant, virtually all of which deserve further scrutiny. Some are clear contract implementation errors, others are likely new, unreported thefts, and yet others capture bridge behaviors that are deserving of considerably more transparency (e.g., unilateral issuance of millions of dollars worth of uncollateralized tokens).
  • $\triangleright$
    Real-time auditing. We show our approach is not only useful for retrospective analysis, but can be used to audit bridge transactions and detect new attacks in real-time. We implement and demonstrate an online audit of the Wormhole bridge and show that it automatically detects every attack we inject. Further, we introduce and implement a proof-of-concept for a new bridge architecture, called announce-then-execute, that incorporates such auditing into the transaction flow itself — thereby preventing unbalanced malicious bridge transactions from ever completing. Our approach treats the most complicated components of bridges (e.g., verifying relayed messages) as black boxes, adds no new attack surface for theft, and requires minimal changes to existing codebases — a benefit which which we demonstrate by implementing a modified version of the Wormhole bridge.
We argue that this approach is powerful both due to its simplicity (in a legitimate financial transaction, the value paid should be equivalent to the value received, less costs) and its independence from the vagaries of smart contract details. Moreover, it represents a straightforward mechanism that can prevent an entire class of existing attacks on crypto bridges.
Figure 1: Cross-chain token bridging and the different steps attackers can exploit to withdraw unbacked deposits.
2 Background
In this section we first give a brief introduction to smart contract blockchains and cross-chain bridges. We then describe how bridges, under attack, collapse in practice.
2.1 Smart-contract Blockchains
Modern DeFi protocols (e.g., decentralized exchanges and lending protocols) are built on top of “smart-contract” blockchains, like Ethereum. At their core, these chains, like the original Bitcoin blockchain, are distributed ledgers that manage accounts (public keys) and their balances (in native tokens like Ethereum’s ETH). Users interact with these chains by signing and broadcasting (to the distributed nodes that make up the blockchain network) transactions that, for example, transfer funds from their account (using the corresponding account private key) to another user’s account.
Ethereum, and the smart-contract chains it inspired since its release in 2015, differ from Bitcoin by extending the “simple” distributed ledger with a smart contract execution layer. A smart contract on Ethereum is an internal account—and, like a normal, externally owned account (EOA), it has a balance—that has associated code (EVM bytecode), which implements the smart contract’s program logic, and storage, which persists the program’s state across executions. Users interact with (i.e., execute) smart contracts much like they do when transferring funds from their account: they sign a transaction that encodes the smart contract to call (i.e., the contract address), the particular function to execute, and the arguments to call the function with. Instead of simply transferring funds from the user’s account, then, executing such a smart-contract call transaction amounts to executing the smart contract bytecode—and any smart contracts the contract itself calls.
Figure 2: Simplified USDC ERC-20 Token Contract.
2.2 ERC-20 Tokens
One of the immediate applications of smart contracts—and to date still one of most popular—is to create custom tokens (or coins). To launch a new token, an organization no longer needs to launch a new blockchain; they can instead deploy a new contract that implements the ERC-20 token standard interface on a chain like Ethereum and take advantage of existing on-chain infrastructure like decentralized exchanges that make it easy to, for example, trade one kind of token for another—both native tokens (e.g., ETH) and other tokens.222 Most EVM chains, i.e., chains that use Ethereum’s execution layer, follow Ethereum standards like ERC-20. Most non-EVM chains like Solana (which have different execution models) have similar standards (e.g., SPL tokens in Solana’s case).
Figure 2 shows a simplified variant of one such token contract—the USDC stablecoin contract. This contract tracks how many USDC tokens an account has and governs the spending of these tokens (much like a bank governs bank notes). For example, the contract’s mint function lets Circle (the company that owns the USDC contract) mint new tokens into a user’s account—e.g., after receiving the corresponding payment from the user off-chain (in US dollars, as USDC tokens are pegged to the US dollar). The contract exposes the ERC-20 interface that lets users (and smart contracts) transfer tokens from their account by simply calling functions like safeTransferFrom, and, in turn, use the tokens in any DeFi protocol (e.g., lending the USDC to different markets, exchanging USDC for ETH, buying NFTs using the USDC, etc.). Finally, the contract’s burn function “burns” a token out of circulation—and emits an event (Figure 3) that Circle’s off-chain code looks for before allowing a user to withdraw the corresponding USD fiat off-chain.
Figure 3: Event Emitted by an ERC-20 Token (USDC).
2.3 Cross-Chain Token Bridges
While smart contracts make it easy to launch new tokens without spinning up new chains, there is no real shortage of new blockchains being deployed (almost weekly). Indeed, the modern blockchain ecosystem is a many-chain ecosystem. Blockchains like Avalanche, Base, and Solana have different design points—from cheaper “gas” execution costs, to higher throughput, lower latency, and different permission models—that make them better suited for different classes of application. And, this has resulted in applications that span many chains and cross-chain infrastructure that ultimately (try to) allow users to, for example, buy Dogwifhat NFTs on Solana using USDC tokens on Ethereum.
Core to these applications and infrastructure—and the focus of our work—is the cross-chain token bridge. At a high level, a cross-chain token bridge makes it possible for a user to “transfer” their tokens from one chain (e.g., their USDC on Ethereum) to another chain (e.g., Solana) and then use the transferred tokens on the destination chain (e.g., on a Solana exchange trading USDC and Dogwifhat). [^3] While some cross-chain bridge do let users transfer one kind of token (e.g., USDC on ETherem) for a completely different kind of token on the destination chain (e.g., Dogwif on Solana), these bridges essentially fuse the cross-chain token bridge with an exchange (or swap). We focus on token bridges not only because they are fundamental to other kinds of bridges, but also because they have higher volume, more liquidity, and more attacks. Since smart contracts cannot make network requests or otherwise access sate outside their own storage, a cross-chain token bridge consists of smart contracts on both the source and destination chains, and off-chain infrastructure that serves to relay the “transfer” call across the two contracts.
As Figure 1 shows, a typical cross-chain token transfer, i.e., a bridge transaction, consists three phases:
1. Deposit (on source chain)
To initiate a cross-chain token transfer, the user first calls the bridge’s contract on the source chain:
deposit(address token, uint256 val, address to) {
address from = msg.sender; // user
address to = address(this); // bridge contract
// ... validate the ERC-20 token contract
token.safeTransferFrom(from, to, val);
emit Deposited(id++, token, from, to, val);
This contract function—a simplified version of the Qubit bridge deposit function—processes the user’s deposit by validating the transfer request, e.g., against a list supported tokens, and then transferring the user’s ERC-20 tokens to the bridge contract—recall contracts are accounts with balances. Then, the function emits an event recording the user’s deposit details (including the deposit ID, the ERC-20 token contract address, the recipient on the destination chain, and value).
2. Off-chain relay
The emitted event is observed by an off-chain relaying component (step 4), which constantly monitors the source blockchain. The relaying component first verifies the authenticity of the deposit event. If the event is authentic, the relaying component then produces a signed receipt endorsing the deposit (step 5) and, typically, stores the receipt off-chain (step 6). Finally, this signed receipt is sent to the bridge’s withdrawal contract on the destination blockchain (step 7). Who submits the receipt varies across bridges—some submit the receipt on the user’s behalf (in these cases, the signed receipt is simply a signed contract-call transaction), while others give users (and anyone willing to pay gas) the signed receipt and they, in turn, submit the receipt to the withdrawal contract to complete the transfer.
3. Withdraw (on destination chain)
The withdrawal contract on the destination chain, first processes the withdraw request by verifying the receipt and transferring the tokens to the recipient (step 8). In (the simplified) Chainswap’s withdraw case, for example, users call:
withdraw(uint256 id, address token, address to, uint256 val, Signature[] sigs) {
_chargeFee();
// verify receipt
require(received[id][to] == 0, ’withdrawn’);
for(uint i=0; i < sigs.length; i++) {
verify_receipt(sigs[i], id, token, to, val);
received[id][to] = val; // mark as withdrawn
token.safeTransferFrom(address(this), to, val);
emit Withdraw(id, token, to, val);
This contract function first charges the caller a fee, then verifies the deposit details against receipt—both that the deposit was not already withdrawn and that the receipt signatures are valid—and finally transfers the tokens to the intended recipient. We consider a cross-chain transaction complete when the asset is released to the recipient (step 9) as such.
We expect every cross-chain bridge transaction to uphold the balance invariant: the value (and kind) of the tokens withdrawn—the outflow—should equal the value (and kind) of the tokens deposited—the inflow—minus the charged fees. In practice, they don’t.
2.4 How Bridges Collapse
In practice, attackers exploited bugs in all three components—the deposit contract, the relayer, and the withdraw contract—and stole signing keys to siphon hundreds of thousands of dollars. Figure 1 highlights the precise steps in the cross-chain token transfer that attackers have historically exploited, including:
  • $\triangleright$
    Bugs in the deposit contract. In step 2, the bridge contract verifies that it has received the correct amount of assets before emitting an event. Bugs in this verification logic could allow an attacker to deposit a smaller amount of assets than what is recorded by the bridge in the event (step 3). For example, the Qubit’s deposit(..) function (see above) did not properly validate the token address; this allowed an attacker to pass 0 for the token address, which meant the contract function did not actually transfer any funds from the attacker’s account but still emitted a Deposit event which allowed the attacker to withdraw actual tokens on the destination chain [25].
  • $\triangleright$
    Bugs in the off-chain deposit verification. In step 5, the relaying component verifies the authenticity of the deposit event emitted by the bridge contract, including whether the event is emitted by the bridge’s designated contract. Bridges that do not correctly verify deposits would allow attackers to withdraw assets that are never deposited.
  • $\triangleright$
    Stolen relayer (or submitter) keys. In step 5, the relaying component signs the deposit receipts which are then submitted to the withdrawal contract as evidence of a valid deposit. If the relayer key is compromised (e.g., was the case with the Ronin bridge [28]) the attacker can forge a valid receipt and then withdraw assets that were never deposited by calling withdraw with the forged receipt.
    The same is true for bridges that submit receipts on behalf of users—and essentially restrict the withdraw callers to privileged submitter accounts. The bridge submitter keys (step 8) have similarly been compromised (e.g., as was the case with AnySwap [4]) and used to withdraw unbacked deposits.
  • $\triangleright$
    Bugs in the withdraw verification. In step 8, the bridge contract verifies that the messages are signed by the relaying component and have not been replayed. Bugs in this verification logic (e.g., as we saw with Wormhole [3]) have allowed attackers to supply “valid” payloads that were not signed by the relaying component and replay withdrawal requests with valid deposits receipts that have already been withdrawn.
In this paper, we assume an attacker that exploits any of the aforementioned components or otherwise controls the relayer (or submitter) keys. In the next section, we show that this attacker model and our simple balance invariant checking captures the largest attacks on cross-chain bridges that have happened in the past. In Sections 5 we show that monitoring withdrawals and deposits on the source and destination chains can be used to detect similar attacks in the future. Finally, in Section 6 we show a describe a announce-then-execute bridge design that enforces this invariant to prevent attacks before they happen.
We note that while our threat model captures a wide variety of vulnerabilities, it is not exhaustive. As with any detection system, an attack that violates one of our assumptions (e.g., avoids violating the balance invariant by transferring funds off-bridge or not having a withdrawal, subverts the transaction data used to validate the invariant, etc.) might succeed. We similarly consider other smart contract bugs (beyond bugs in deposit and withdrawal functions) and account key compromises out of scope—and instead focus on the cross-chain bridging aspects which are relatively less well understood. As we show later, our model captures the largest attacks on cross-chain bridges that have happened in the past and systems can use it to prevent similar attacks in the future.
3 Approach
The core hypothesis of our work is that value should be conserved within cross-chain transactions. That is, that the value of the asset inflow in such a transaction (i.e., the deposit) should equal the value of the asset outflow (i.e., the withdrawal). In token transfer bridges, this invariant corresponds to a balancing of the inflow tokens and the outflow tokens (less any fees or transaction costs incurred by the bridge itself). When this balance invariant does not hold it allows a range of opportunities for fraud, all of which involve greater outflows (withdrawals) than inflow (deposits). Figure 4 illustrates such an outcome by graphing the aggregate difference between bridge inflow (via Ethereum) and outflow (via Solana) leading up to and during the February 2022 attack on the Wormhole bridge. The net difference is consistently near zero, with only short positive deviations (representing delayed withdrawals) until the attack in January, at which point there are significant withdrawals without matching deposits—producing a large negative difference.
Testing this invariant on a per transaction basis, is straightforward in principle, but since bridge transaction formats are not standardized, it requires a range of per-bridge and per-chain parsing in practice. Our methodology for normalizing this information focuses on two key pieces of information: a) identifying each bridge transaction—a composite of a deposit (inflow) transaction on a source blockchain and a withdrawal (outflow) transaction on another—and b) identifying the value transferred in each such bridge transaction.
Figure 4: Total Inflow (ETH) - Total Outflow (SOL) Over Time: Wormhole Attack in Feb 2022.
3.1 Identifying Bridge Transactions
For almost all bridges, identifying their component (per-chain) inflow and outflow transactions is straightforward — bridges typically use explicit events on each chain to signal if a given transaction is a deposit or withdrawal. [^4] One key exception to this rule is the Wormhole bridge which, until late 2023, did not emit specific events when executing a withdrawal transaction. In this case, we infer that a withdrawal took place by looking for transactions that invoke functions designated for performing withdrawal operations (e.g., completeTransfer). It also appears to be widely understood today that emitting explicit events is a best practice.
Pairing these component transactions (i.e., matching deposits on one chain to withdrawals on other) can be performed in several different ways. The easiest, and most common, is via a unique transaction identifier. Such ids are typically generated by the bridge on the source chain during a deposit transaction and then copied into the withdrawal transaction on the destination chain. [^5] These unique ids are commonly simple global variables incremented with each new transaction. However, instead of an explicit id, some bridges use a hash of the deposit transaction for the same purpose (e.g., for Anyswap, each withdraw transaction will include the hash of its corresponding deposit transaction). [^6] We note that to prevent potential replay attacks, bridge implementers should also include an id that unique identifies individual deposit events within a transactions. Sadly, this is often not the case. Finally, in a handful of cases there are no “inband” identifiers that can be used to associate transactions. We believe this is a poor design choice that is fundamentally in conflict with transparency. However, even in these cases, for the purpose of our analysis we have been able to pair transactions using explicit query APIs provided by the affected bridges services. [^7] For example, when a bridge transaction on the Poly Network bridge includes a withdrawal from Curve (a kind of liquidity pool) or when a bridge transaction on the Binance Token Hub includes a Binance Smart Chain (BSC) withdrawal, it is not possible to identify the partner deposit from blockchain data alone and we must make use of “out of band” data available through their respective bridge query APIs.
At the end of this process, we identify a comprehensive collection of “bridge transactions” (a pair of transactions from two different blockchains that were used by the bridge to transfer of value across them). While the vast majority of bridge transactions are pairs of deposit and withdrawal transactions, a handful of bridge transactions are either withdrawal-only (e.g., in the case of attacks) or deposit-only (e.g., if the user chooses to delay their withdrawal).
3.2 Identifying the Value Transferred
Many bridge transaction event formats explicitly identify the amount and kind of tokens transferred. However, some bridges do not emit this information and others can be unreliable. [^8] For example, as mentioned earlier, pre-2024 Wormhole does not emit an event for withdrawal transactions. Other examples include the Meter bridge and Qubit bridge, which attempted to verify the number of tokens received but had exploitable bugs. In such cases, we can frequently make use of the ERC-20 “Transfer” event (see Figure 3 for an example) that is emitted when the contract transfers the tokens from the user’s account to the bridge’s account. This event contains the number of tokens transferred as well as the sender and recipient addresses, which we use to identify the number of tokens transferred to or from the bridge. Because the Transfer event is adjacent to its associated bridge-generated event, it is easy to identify and thus establish the number of tokens transferred. [^9] As a sanity check, we also require that the sender and recipient addresses in the Transfer event are consistent with the bridge’s defined behavior: for withdrawal transactions, we require that the sender is either a mint address (i.e., all-zero address) or a bridge-controlled address, while for deposit transactions, we require that the recipient is either a burn address (i.e., all-zero address) or a bridge-controlled address. In a few implementations, multiple related Transfer events can be emitted at once, and in these cases we have manually inspected their contracts and constructed implementation-specific logic to account for this behavior. Another special case is caused by so-called “reflection tokens” in which the number of tokens logged in the event (the value field in Figure 3) is dynamically adjusted based on a combination of the intended number and the total token supply. For such cases, we either rely on bridge events which capture the intended number of tokens transferred or implement token-specific logic to recompute the value accordingly. Finally, native tokens have no Transfer event (since they are not ERC-20 tokens), but thus far we have either been able to recover the number of tokens transferred from bridge events or so-called “internal transactions”.
To summarize, while it is certainly possible to create a bridge transaction protocol that records insufficient data to match deposits and withdrawals, or for which the number of tokens transferred might be ambiguous, our empirical experience analyzing 11 bridges and 20 blockchains is that such reconstruction has always been possible.
3.3 Checking the Balance Invariant
Once we have identified both sides of the bridge transaction and the number of tokens transferred by the bridge, we can verify if the balance invariant holds: for each bridge transaction (including those that are withdrawal-only), does the number of tokens transferred from the bridge—the withdrawal amount—match the number of tokens received by the bridge—the deposit amount?
However, a key complication is that bridges can charge fees and, while these fees can sometimes be paid “out of band”, it is not uncommon for them to be subtracted from the tokens received on deposit. To account for this behavior (i.e., outflow = inflow - costs), we must be able to determine such costs on a per-transaction basis. In most cases, fees can be accounted for in a straightforward manner: they either are made explicit in bridge-generated events (and can thus be accounted for directly) or can be calculated based on either published fee schedules or inferred fee schedules (i.e., since all transactions are typically subject to the same fixed or percentage fees).
In some cases, the precise value of fees may be difficult to determine retrospectively because they depend on some external contemporaneous value not recorded in the transaction (e.g., fees valued in US dollars implicitly depend on the exchange rate of a given token at that time). While such ambiguity could be further minimized with additional data, in our work we manage this issue by defaulting to a simple rule that the number of tokens withdrawn should not exceed the number deposited.
Blockchain Bridges that Operate on the Chain
Arbitrum Anyswap, Poly Network, Wormhole
Avalanche Anyswap, Poly Network, Meter, Nomad, Wormhole
Binance Beacon Binance Token Hub
Binance Smart Anyswap, Binance Token Hub, Chainswap, Harmony, Meter, Omni, Poly Network, Qubit, Wormhole
Celo Anyswap, Poly Network, Wormhole
ETH Anyswap, Chainswap, Harmony, HECO, Meter, Nomad, Omni, Poly Network, Qubit, Ronin, Wormhole
EVMOS Anyswap, Nomad, Omni
Fantom Anyswap, Poly Network, Wormhole
Gnosis Omni, Poly Network
Harmony Anyswap, Harmony Bridge, Poly Network
HECO Anyswap, Chainswap, HECO, Poly Network
Meter Meter
Metis Anyswap, Poly Network
Milkomeda Nomad
Moonbeam Anyswap, Meter, Nomad, Wormhole
Moonriver Anyswap, Meter
OKT Chain Anyswap, Chainswap, Poly Network
Optimism Anyswap, Poly Network, Wormhole
Polygon Anyswap, Chainswap, Poly Network, Wormhole
Ronin Ronin
Solana Wormhole
Table 1: The blockchains CrossChecked supports and the cross-chain bridges that operate on them. While a particular attack on a bridge involves two chains, we collect deposit and withdrawal transactions for a bridge on all chains that the bridge supports.
Name Date Loss (USD) Analyzed Reported New Test Error Suspicious
Ronin Mar 2022 $624.0$m 3.0m 2 - - - -
PolyNetwork/2021 Aug 2021 $611.0$m 292k 18 - - 1 -
BSC Token Hub Oct 2022 $587.0$m 2.0m 2 - - - -
Wormhole Jan 2022 $360.0$m 642k 1 - - 2 -
Nomad Aug 2022 $152.0$m 37k 962 - - - -
Harmony Jun 2022 $100.0$m 336k 15 - - - 43
HECO Nov 2023 $86.0$m 23k 8 - - - 73
Qubit Jan 2022 $80.0$m 260 16 - 114 - -
Anyswap Jul 2021 $7.9$m 3.4m 4 24 - 806 7
PolyNetwork/2023 Jun 2023 $4.4$m 290k 136 - - 1 27
Chainswap Jul 2021 $4.4$m 53k 1136 - 15 4 -
Meter Feb 2022 $4.3$m 14k 5 - - - -
Total $2.6$b 10.1m 2,305 24 129 814 150
Table 2: List of top attacks on cross-chain bridges in the retrospective analysis, ordered by amount stolen. CrossChecked analyzed over 10m bridge transactions (20m component deposit and withdrawal transactions) and identified all bridge transactions previously reported as having been associated with the attacks (Reported). It also identified bridge transactions that violated the invariant that were a previously unidentified attack (New), test transactions (Test), transactions reflecting bugs in implementations or use (Error), and suspicious transactions that employ manual signing (Suspicious). CrossChecked identified slightly more transactions than were reported by the Nomad [20] and Meter [22] bridges for their attacks (+2 for Nomad, +1 for Meter). The Chainswap attack only has reports of the malicious deposit address [18], which matches the one identified by CrossChecked.
4 Retrospective Analysis
The basic reasoning motivating the balance invariant is simple: that legitimate bridge transactions should conserve value. However, this assumes a particular model for how bridges are used and operated which may or may not hold in practice. To validate our hypothesis, we applied the balance invariant analysis retrospectively across a large set of past transactions which, while primarily benign, contain the largest known bridge attacks during our period of study. Our goal is both to show that all real attacks are identified (detection), but also that the invariant does not alert on large numbers of benign transactions (bridge compatibility). Later, in Sections 5 and 6, we will describe systems for live bridge monitoring as a third party and a new implementation for preventing unbalanced transactions from being committed.
4.1 Data Set
To perform a retrospective analysis we need to data for which (at least some) of the ground truth is known. In this section we describe how we chose the bridge attacks we study, the blockchains and smart contracts involved, and how we collected the historical transaction data.
Attack Selection.
We compiled a comprehensive list of attacks on cross-chain bridges that occurred between January 2021 and December 2023. These were sourced both from academic papers surveying cross-chain bridge attacks [14, 36, 38] as well as industry blog posts that collect and characterize attacks over time [1, 27, 6, 32, 31]. As this paper focuses on end-to-end auditing of bridge transactions, we filter out attacks involving blockchains lacking smart contracts such as Bitcoin (e.g., the pNetwork attack in 2021), attacks on swap bridges (e.g., the Thor bridge attacks), attacks that do not involve individual bridge transactions (e.g., the evoDefi attack), and attacks that withdraw funds through other means (e.g., direct transfer from a vault that stores assets for a bridge such as the Multichain 2023 attack). Moreover, given the large number of attacks during this period, we focus on major attacks with claimed losses greater than \$1 million USD and exclude the remaining (the sum of losses from these excluded attacks are a small fraction of those in our scope). This process yielded 12 attacks on 11 distinct bridges. We note that this list includes the top five attacks on cross-chain bridges in history, which collectively resulted in more than \$2.6 billion USD in claimed losses.
Blockchain Selection.
Validating bridge transactions requires access to transaction data from both the source and destination chains. We support every blockchain involved in the bridge transactions associated with the 12 attacks, either as the source (or claimed source) or destination chain. As a result, we support a total of 21 blockchains that together cover a broad range of designs, including many of the most popular such as Ethereum, Binance Smart Chain, and Solana. [^10] Details about some deposit transactions on PolyNetwork (deposits that are used for liquidity purposes) are only available through PolyNetwork’s API. We support querying PolyNetwork’s API, but do not list it as a supported blockchain. Table 1 lists the blockchains and the bridges in our retrospective study that operate on them.
Smart Contract Selection.
For each bridge, we comprehensively collected the results of all versions of its bridging smart contracts on every blockchain we considered. In particular, we collected deposit and withdrawal transactions created by every contract the bridge deployed on every blockchain we support (typically many more blockchains than the ones involved in a bridge attack) as well as all versions of the smart contract implementation for the bridge (bridges evolve their implementations over time, such as in response to an attack). Since the transactions we collected from this set of smart contracts are much broader than those just involved in the top 12 attacks we consider, including them further reinforces our findings that the alerting workload is very small (Section 4.2).
Figure 5: The lifetime of the bridges in our retrospective study. Lines start with the bridge’s first valid transaction and end with the last valid transaction in our data, corresponding to the bridge’s closure or November 2023 if the bridge was still operating at the end of our data set. Diamonds indicate the dates of attack.
Data Collection.
Collecting smart contract transaction data—the verbose record of what each contract did—across many chains constitutes the most time-intensive aspect of performing a retrospective analysis.
For each bridge, we collected deposit and withdrawal transactions for each of the blockchains it operates on (for the 21 chains we support) using commercial RPC services which charge for queries. We primarily used five commercial services, as no single service supports all of the blockchains and functionality we need. [^11] These five RPC services are: GetBlock, QuickNode, ChainStack, GoldRush, and Ankr. For each bridge and blockchain combination, by default we collected all of the historical transactions generated by its smart contracts from the genesis of its deployment to the end of November 2023 or the end of its life, whichever came first.
The two exceptions to this rule are the Binance bridge, for which we limited data collection to a year (six months prior to its attack and six months after) because of the sheer volume of transactions over its lifetime (30 million transactions) and the Harmony bridge, whose contract was reused by another bridge (LayerZero) after its attack (since LayerZero was not attacked itself, we only consider transactions that were directly associated with Harmony).
Figure 5 illustrates these bridge lifetimes with each line starting at a given bridge’s first transaction and ending with its last in our data set. Diamonds indicate the dates of individual attacks, emphasizing that many of the bridges closed shortly after their attacks.
4.2 Analysis
We developed a tool called CrossChecked that pairs deposit and withdrawal transactions from blockchains into bridge transactions, and applies the balance invariant and consistency checks on them. Using the transactions we collected for the 12 attacks across 11 bridges and 21 chains, CrossChecked analyzed over 10m bridge transactions, identifying more than 2.3k bridge transactions associated with the attacks and 1.1k more that violated the balance invariant.
Table 2 summarizes our results. For each attack, it shows the bridge involved, the date of the attack, the claimed loss in USD, the number of transactions CrossChecked analyzed, and the number and kinds of transactions that violated the balance invariant. Below we discuss these two categories of transactions in more detail. For further context on each of the attacks, Appendix A provides a descriptive account summarizing how the attacks transpired.
Looking forward to Sections 5 and 6, altogether CrossChecked identified 3,423 bridge transactions (0.03%) that violated the balance invariant out of more than 10m analyzed. If the invariant is used by a third-party auditing or protection system, we note that such an alert workload has a negligible overhead for manual inspection, typically raising no more than one alert (or one batch of alerts) every few weeks.
4.2.1 Reported
For each of the 12 significant cross-chain attacks, Section 4.1 describes how we gathered the historical transactions that correspond to the attacks using external sources. We use these identified transactions as ground truth for evaluating the ability of CrossChecked to identify attack transactions. As shown in Table 2, when processing the transaction histories of the chains involved, CrossChecked successfully identified all 2,305 bridge transactions on the source and destination chains associated with the attacks.
Since we designed CrossChecked to specifically identify such attacks, these results may not be surprising. However, they are useful for confirming that the approach of checking a simple, well-defined invariant works well. Moreover, the approach works well across a variety of models, including bridges that specify fees in a fiat currency, tokens that use a reflection mechanism, etc.
4.2.2 Other Violating Transactions
Equally compelling is the extent to which other, non-attack transactions violate the balance invariants. If the attack transactions are dominated by many false positives, then the approach becomes less effective. As shown in Table 2, CrossChecked finds significant fewer (1,117 compared to 2,305) bridge transactions that were not previously identified as attacks. Given the nature of these additional transactions, though, they do not undermine the effectiveness of the invariant approach. By violating the invariant something highly unusual is taking place. As a result, we argue that such transactions should be flagged for further scrutiny and perhaps even blocked from executing (particularly transactions in the “new” category and the large transactions in the “suspicious” category below).
To ensure that we have not missed benign explanations for a violation, we manually inspected violating bridge transactions in at least of one of the following ways: (1) using blockchain explorers to verify that the funds have been created by the withdrawal (and if so, whether they have been moved); (2) searching online for any additional information about the transaction and addresses involved; (3) checking that if claimed deposit exists; and (4) examining unredeemed deposits on the source chain that potentially could have been used to back the withdrawal (e.g., because the smart contract implementation changed between the deposit and withdrawal). If we can manually reconcile a bridge transaction, we consider it benign and do not consider it further.
We group the remaining bridge transactions that violate the invariant into four categories, which we describe below. For reference, we also list some of these bridge transactions in Table 3 in the appendix to provide specific examples with more detail.
New. We believe that we have identified previously unreported bridge transactions involved in two new, unreported attacks on Anyswap. The first group of three transactions executed once Anyswap reopened after the attack but before Anyswap patched its smart contract (Anyswap transactions #1–3 in Table 3 in the appendix). These transactions were three days later and involved different deposit and withdrawal addresses than the July 10, 2021 attack. The second group of 21 transactions on November 18, 2021, were all withdrawing on Avalanche. These transactions either referenced deposits that had already been redeemed months previously (Anyswap #6 in Table 3) or referenced non-existing deposits (Anyswap #7 in Table 3). Manually inspecting the smart contract used, the attackers were exploiting a bug in Anyswap’s implementation (the access control code was commented out and thus ineffective) and one of the addresses was labeled as “KyberSwap Exploiter”.
Tests. Two bridges had test transactions that violated the invariant. Chainswap had 15 transactions that withdrew tokens labeled as test tokens (e.g., tokens labelled as “testtoken” or “startertoken”). Similarly, Qubit had 114 transactions that minted test tokens (e.g., “xTST”) without backing deposits. While all these transactions did not have corresponding deposits, we surmise that they were likely benign given the tokens involved (tokens that have no real value).
Error. Four bridges had transactions that suggest bugs in either their implementation or their invocation.
PolyNetwork/2021 had one bridge transaction where the withdrawal amount matched the deposit amount, but the withdrawal moved the funds to the wrong destination chain (CrossChecked flagged the mismatch in destination chain specified in the deposit and withdrawal transactions). Anyswap had two groups of bridge transactions in this category. The first consists of 800 withdrawal transactions that pointed to just a few deposits. Manually sampling a few, the deposit transaction hash was not set correctly in the withdrawal transactions (we found their matching deposits), suggesting a program error. The second Anyswap group had four double-spend transactions (referencing the same deposit) that are also likely program errors: the deposit and withdrawal amounts are the same, and the withdrawals occurred within a few hours of each other. Wormhole had two bridge transactions that violated the invariant: 0.5 wrapped Solana on Polygon $\rightarrow$ 650 USDC on Solana, and 4.3k MATIC on Polygon $\rightarrow$ 2.3k on Avalanche. The small amounts and close proximity of the dates of the transactions suggest they are also likely errors.
Finally, three bridges had transactions that had no apparent effect, suggesting invocation errors or undocumented behaviors. PolyNetwork/2023 had one bridge transaction, and Chainswap had four, where the deposits were non-zero but the withdrawal amount was zero. And Anyswap had two withdrawals referencing deposits that did not specify a recipient, yet the deposit amounts covered the withdrawals. While technically not balance invariant violations, CrossChecked flagged them because of their unusual circumstances.
Suspicious. The final category of bridge transactions suggests the manual, intentional use of private keys in signing transactions that effectively bypass verification — precisely the kind of transactions that warrant auditing.
Many of these cases involved highly suspicious unbacked bridge transactions involving very large withdrawals without corresponding deposits. For example, Anyswap had seven transactions totaling more than \$1.5m pointing to non-existent or already-redeemed deposits (e.g., Anyswap #4 and #5 in Table 3 in the appendix). And PolyNetwork/2023 had 27 withdrawals referencing a non-existant deposit address totaling more than \$20m (PolyNetwork/2023 #1 and #2 in Table 3).
The HECO bridge had 73 unusual bridge transactions. One was a very suspicious unbacked bridge transaction minting \$5m of USDT on the HECO chain (HECO #1 in Table 3). The remaining 72, totaling over \$36m, involved withdrawals to an address labeled “HECO recovery”. The label suggests benign intent such as rescuing funds trapped in the bridge, but the activity is also consistent with a rug pull.
Other cases suggest seemingly careless operational practices. In particular, Harmony had 43 bridge transactions that violated the invariant in a variety of different ways. Eight double-spending bridge transactions (e.g., Harmony #1 in Table 3 in the Appendix) used the same deposit to release tokens twice (though only resulting in a profit of a few hundred USD). Thirty-two had indecipherable data: it was not possible to decode the deposit function name, function input, and some events (thus preventing verification of the deposit). One bridge transaction minted piggybankone tokens on Harmony chain referencing a non-existing deposit. And two had withdrawal amounts that were smaller than the deposits, perhaps caused by undocumented fees or errors. Considering the specific mechanism used by the Harmony bridge, where a private key was signing transactions that it should not have, and the fact that some double-spending transactions had amounts different than what was deposited, these incidents suggest that Harmony had issues operating securely and correctly.
5 Live Auditing
Section 4 showed the effectiveness of applying the balance invariant on restropective data to accurately identify past bridge attacks. Incorporating the logic used in the retrospective analysis, we implemented a real-time auditing system that uses the balance invariant to monitor ongoing transactions on a bridge.
Since most of the bridges in our retrospective analysis shut down at the time of writing (Figure 5), we focus on the Wormhole bridge. Wormhole is still operational, supports a wide variety of blockchains, and is among the most popular bridges in terms of transaction activity [36]. This live system monitors thousands of withdrawal transactions per day across ten blockchains, which together account for 60% of all withdrawal transactions on Wormhole. Extending the system to support other bridges is straightforward.
5.1 System Overview
Figure 6 depicts the architecture of our live auditing system. The system consists of three main components. The Blockchain Monitor tracks the blockchain network and retrieves the latest deposit and withdrawal transactions. The Auditor checks extracted transactions for invariant violations. The Database component stores all the transactions and auditing results.
Figure 6: Live auditing system which consists of three components.
Blockchain Monitor.
The Blockchain Monitor obtains deposit and withdrawal transactions by periodically retrieving the latest finalized blocks. The Monitor uses the same commercial RPC services for collecting Wormhole deposits and withdrawals as we used for the retrospective analysis (Section 4.1). [^12] A fully operational deployment would manage nodes participating in each of the blockchains to obtain transactions in block data directly.
The Monitor only retrieves finalized blocks since unfinalized blocks may be reverted or reordered. Since different blockchains have different finality times, we use the timestamp of the latest finalized block of Ethereum as the synchronization time since Ethereum has the slowest finality time (around 12 minutes) [12]. After retrieving the blocks, the Monitor extracts deposit and withdrawal transactions for Wormhole and saves them to a local database.
The polling interval is configurable and the Monitor polls chains every 1 minute by default. As a result, the live auditing system is as recent as the polling interval combined with the finality time — a situation similar to bridges due to the polling nature of distributed off-chain communication. We leave it to future work to optimize the delay further.
Auditor.
For every new withdrawal transaction extracted by the Blockchain Monitor, the Auditor attempts to find the corresponding deposit transaction in its local database. In most cases it will find the deposit, but if it cannot find it (e.g., because the withdrawal references a non-existent transaction), the Auditor sends an email alert inviting manual inspection. If it finds a deposit, it examines the deposit and withdrawal transactions, and sends an alert if either the deposit has already been withdrawn (the new withdrawal is double-spending) or the bridge transaction violates the balance invariant.
Database.
The system uses a PostgreSQL database to store all information regarding deposit and withdrawal transactions. This information includes the transaction hash, the blockchain, the amount, and the timestamp. It also stores the auditing results, including whether the deposit has been withdrawn and whether the invariant is violated.
5.2 Deployment
We have deployed our live auditing system for the Wormhole bridge for a month. The system audits the deposit and withdrawal transactions between 10 blockchains, which collectively account for around 60% of all withdrawal transactions on Wormhole. On average the observed throughput is approximately one new transaction every two minutes.
The system currently refreshes every one minute, and sends email alerts if it observes a transaction violating the balance invariant or another auditing property (e.g., the destination chain in the deposit does not match the chain in the withdrawal). In the time that we have had the system operational, it has alerted on multiple transactions that violated the invariant. So far these transactions have proven to be errors in our implementation (e.g., failed to extract the correct amount) or the RPC services we used (e.g., not indexing a transaction).
In addition, while we have not observed any attacks during our monitoring period, we simulated three attack scenarios to confirm that the system alerts when expected. These three transactions represent three types of attacks: (1) a double-spend attack, (2) an unbacked withdrawal attack, and (3) an attack where the deposit and withdrawal amounts do not match. The system successfully alerted on all three attacks.
6 Protecting Bridges
Sections 4 and 5 showed how applying the balance invariant can identify attacks retrospectively and monitor ongoing transactions as an external third party without any changes to existing bridge infrastructure. However, the live monitoring system only alerts on violations of the invariant and does not prevent violating bridge transactions from being processed.
In this section, we describe an approach for extending an existing bridge implementation to prevent malicious transactions from executing. Our goal is to present a proof-of-concept implementation to demonstrate the feasibility of applying the balance invariant in the workflow of bridge transactions as an active defense against malicious transactions. We acknowledge that a number of practical issues remain for a complete operational system, presenting an opportunity for interesting future work.
We start by proposing a new model for the workflow of bridge transactions. Then, using Wormhole as an example, we show how to modify an existing bridge implementation to prevent malicious transactions from being processed with fewer than 30 lines of code changes. Finally, we evaluate the correctness and overhead of this initial implementation.
6.1 The Announce-then-Execute Model
To guide the design of a new model, we first highlight where the attack surfaces exist. As shown in Figure 1, a vulnerability can exist in as early as the second step (verify lock amount) to as late as the next-to-last step (verify relayed transaction). Thus, a natural place to introduce a protection mechanism is just before the final step, which executes the relayed transaction.
Using this insight we propose an announce-then-execute model. Figure 7 depicts the entities and steps involved in the new model, focusing on the destination chain (replacing the right-most box of Figure 1). The new model employs two steps to process a withdrawal transaction: (1) verify and announce the withdrawal transaction on the destination blockchain (steps 9 and 10) and (2) execute the withdrawal transaction (steps 11 and 12). Of particular note, this new model introduces an additional entity, the Approver, which is responsible for approving the relayed transaction before it is executed. The Approver can be a third-party service or a bridge operator, and is responsible for verifying that the announced withdrawal transaction satisfies the invariant by validating the same conditions as the Auditor in the live monitoring system (Section 5). If the relayed transaction violates the invariant, the Approver can reject the transaction to prevent it from being executed.
The announce-then-execute model has a few advantages. First, it explicitly separates the verification and execution of the relayed transaction, thereby providing an opportunity to reject transactions that violate the balance invariant. In comparison, many bridges (e.g., Wormhole) couple the last two steps tightly — the same function call performs both the verification and execution of the relayed transaction. As a result, if the verification step has a bug, it is not possible to stop violating transactions from being executed. Second, the announce-then-execute model allows any external party, that has visibility into both blockchains, to determine if the relayed transaction violates the invariant. A long-standing challenge with bridges is the oracle problem: the destination blockchain does not have visibility into the source blockchain. The announce-then-execute model side-steps the oracle problem by allowing an external party to verify the relayed transaction before it is executed. Moreover, the announce-then-execute model is compatible with existing bridges, as it only requires modification to the last step and treats the relaying infrastructure as well as the verification code as a black box. Lastly, the announce-then-execute model does not introduce any new attack surfaces — even if the approver naively approves all transactions, an attacker still has to exploit other vulnerabilities in the bridge to profit.
Figure 7: Announce-then-execute model for bridges.
6.2 Implementation
Continuing our focus on Wormhole from Section 5, we modified one of Wormhole’s open-source contracts to implement the announce-then-execute model. Wormhole’s current withdrawal function performs two tasks: it verifies the signatures that authorize the withdrawal transaction, and then executes it. We split the withdrawal function into two separate functions: announceWithdraw and approveWithdraw. The announceWithdraw function handles signature verification and announces the withdrawal transaction on the destination blockchain, while the approveWithdraw function executes the withdrawal transaction. This modification required fewer than 30 lines of code changes.
6.3 Evaluation
As a final step we performed a high-level evaluation of the correctness and overhead (in terms of gas usage) of the announce-then-execute implementation. We deployed our bridge implementation on the Binance and Fantom testnets due to the availability of free faucets for acquiring testnet tokens for them. We then executed a series of benign and malicious transactions using Binance as the source chain and Fantom as the destination chain.
Correctness. For evaluating correctness we disabled the security checks in the original implementation (e.g., signature verification) to allow malicious transactions to flow through the contract implementation (e.g., simulating key compromise). We then executed 100 pairs of deposit and withdrawal transactions to validate that the implementation correctly identifies and protects itself against malicious transactions that violate the balance invariant.
All but three of the 100 paired transactions were benign, and used randomly generated values as transaction inputs. The remaining three paired transactions were malicious and were randomly placed in the transaction sequence. The malicious transactions represented the three kinds of bugs underlying the large attacks in Section 4: (1) a bug that allows a user to withdraw more tokens than they deposited, (2) a bug that allows a user to withdraw tokens without depositing any (simulating key compromise), and (3) a bug that allows a user to withdraw twice from the same deposit (double spending).
The bridge implementation correctly executed the benign transactions to completion and rejected the three malicious transactions. The results are the same independent of where the malicious transactions randomly appear in the sequence.
Overhead. The additional steps add overhead to implementing bridges. For this experiment, we measured the gas usage for executing Wormhole’s original implementation (with its security checks) and the gas usage of the announce-then-execute implementation (also with the original security checks plus our added code). On average, our proof-of-concept implementation consumed 110k more gas than the original Wormhole implementation (370k vs. 260k gas) when executing the benign transactions, which translates to a roughly \$1 increase for a typical withdrawal on Ethereum using Wormhole.
We note that our implementation is not optimized for gas usage, and the results indicate that a real deployment of the announce-then-execute model will want to optimize the gas overhead of the additional work (e.g., narrowing to transactions above a threshold amount of funds). As discussed further in Section 8, such optimizations and other practical issues of a real deployment are interesting future work.
7 Related Work
Blockchains are, by their nature, public and thus support direct empirical analysis of past transactions. This has engendered a rich literature quantifying and characterizing a range of quasi-adversarial activities on individual blockchains including arbitrage [19, 24], sandwich attacks [24, 39], frontrunning [5], transaction replays [24], and a range of smart contract vulnerabilities and attacks [23, 10, 26, 37, 33, 9]. While our work focuses on the cross-chain context, a number of the vulnerability classes identified in such work are directly implicated in the attacks we analyze.
In the cross-chain context, there are several different streams of related work. First are the efforts to improve the security and performance of cross-chain bridge designs — particularly in managing cross-chain consensus concerning state changes — using zero-knowledge [34] or committees of validators[13, 15]. Some many-chain blockchain ecosystems (e.g., Avalanche) are extending their chains with built-in bridges (across their different chains). We consider our work orthogonal to these efforts as we focus on unbalanced bridge transactions, independent of the particular security violation that allowed such an outcome to take place. Indeed, some of the compromised bridges did use e.g., set of validators.
Another line of work has reviewed real-world attacks on cross chain bridges[14, 36, 21, 38] — ranging from just a few such events to an analysis of over 30 attacks. Our work has directly benefited from the insight and documentation these authors provide, but our respective focus is distinct. While these efforts have concentrated on identifying the vulnerabilities and mechanisms of attack, our work is agonistic to these details and focuses on the financial side-effects those actions.
Yet a third research direction seeks to automatically discover new vulnerabilities in cross-chain bridges using either machine learning (e.g., ChainSniper [29], which trains models to identify vulnerable smart contracts and Lin et al. [17], which train models to detect fake deposit events) or static analysis (e.g., XGuard [30], which statically analyzes bridge contracts for inconsistent behaviors, and SmartAxe [16], which analyzes the control-flow graph of smart contracts and identifies access control and semantic vulnerabilities).
The papers philosophically closest to ours are those that consider the security properties of cross-chain bridges at a higher-level of abstraction. For example, Belichior et al. [2] build and evaluate a synthetic bridge designed to allow high-level monitoring of bridge behavior — including variations in financial state. In a more empirical context, Huang et al. [11] characterize the transactions of the Stargate bridge, and highlight the correlation between large or unusual trades and individual attacks. Finally, perhaps the closest work to our own is Zhang et al.’s XScope [35] which models real-world bridge bugs using a set of pre-defined rules and applies these empirically to identify several attacks retrospectively. Like us, they focused on invariant patterns than the low-level details of smart contract bugs. That said, the rules considered by XScope are still considerably lower-level and more granular than our balance invariant as they are seeking to identify the cause of the detected attacks that have taken place. While two efforts are complementary, we believe our work is simpler to understand and implement, more clearly robust (we have tested against a far wider range of bridges and blockchains) And, as a result, is likely more attractive for inline deployment.
Finally, a range of blockchain companies advertise tools that actively monitor cross-chain bridge transactions for anomalies. These include Hexagate, Hypernative, PeckShield, Slowmist, Certik and ChainAegis (among others), all of which claim various levels of automated alerting for suspicious transactions. However, without clear information on how these systems operate, or empirical results on their efficacy, it is hard to place our work in context beyond our shared goals.
8 Discussion and Future Work
This paper is a “first cut” at identifying, validating and implementing a generic mode of theft protection into cross-chain transactions. We show that a simple balance invariant is sufficient to detect and prevent most bridge attacks. Our effort also highlights the need for better transparency and auditing in the bridge ecosystem, and for new bridge designs. Indeed, implementing and deploying a cross-chain token bridge with an in-line balance invariant monitor is an open challenge.
As mentioned in Section 2, in this paper we focus on cross-chain token bridges and thus our approach does not directly apply to bridges that have built-in exchanges and swaps. Still, similar invariant-checking techniques likely extend to such cross-chain exchanges and similar protocols (e.g., Automated Market Makers which trade between a range of cryptocurrencies in decentralized exchanges). Such extensions will require incorporating contemporaneous price oracle data into our framework and reasoning about the variation therein.
The larger opportunities opened up by this work are more architectural. Today, when a crypto platform declares that their system has been audited, they typically are referring to a third-party that inspects their smart contracts for common logic errors or deficiencies in their process. Any validation of overall financial safety is inherently entwined with the totality of the implementation. However, such code-oriented audits are inherently challenging and ill-suited for overarching questions of financial risk. They must evaluate all contracts and bridge code that directly or indirectly have access to keys controlling assets and then must ensure that no combination of inputs or invocations might lead to a loss (itself, an outcome that is rarely well defined a priori). Unsurprisingly, multiple bridges in our study had been previously audited by third parties before their attacks.
Our work suggests that by separating overall financial safety invariants from the intricate details of every given contract or trade, we can dramatically simplify the defender’s job and their ability to reason about risk. Similar to so-called “circuit breakers” implemented in traditional securities exchanges, or the atomic repayment guarantees embedded in single-blockchain “flash loan” services, the ability to script overarching financial constraints independent of transaction details can be extremely powerful. For example, one might extend this work beyond a simple cross-chain balance invariant to enforce limits on liquidity losses or to place limits on collateralization risk for tokens backed by stablecoins. We hypothesize that by centralizing the specification and enforcement of overall financial constraints in one place, they will be far easier to audit and reason about.
Open Science Policy
We will make all of our data, results, and code public upon publication. In addition, we have already included some of the most important results in the appendix.
Ethics
We believe our work, which deals with public data, no identified individuals and a simple means for identifying attacks on crypto token transfer bridges (and potentially preventing such attacks in the future), has very low ethical risk and significant upside. However, in the interest of due diligence, we provide a more thorough justification of this position following the Menlo Report principles as indicated by the USENIX Security call.
Considering the Menlo report principle of “Respect for Persons”, we are not dealing with, nor do we attempt to analyze, data at the granularity of individual persons. We strictly use public data and to the extent that we identify outcomes (e.g., attacks) they are at the granularity of bridges and blockchains, not individuals. Similarly, with respect to the principle of “Justice”, we have no reason to believe that our analysis discriminates at the level of persons or distinct groups of persons — all users of bridges should benefit equally (excepting criminals). On the topic of “Respect for Law and Public Interest”, we are engaging with existing public data, we have documented our methods clearly and we are careful to limit our factual statements to those for which we have evidence. As well, we have attempted to disclose significant suspicious bridge transactions to appropriate bridge operators (yet with limited success as most of the bridges have ceased operations).
Finally, on the topic of “Beneficence”, we identify the three sets of stakeholders in our work. These include the past and (potential) future victims whose investments, both direct and indirect, via token transfer bridges have been undermined by large-scale thefts from these services. In addition, other stakeholders include bridge operators and, finally, the criminal actors who derive income from vulnerabilities in bridges. We argue that our retrospective analysis creates no direct new harm for any parties — these attacks have already happened. For bridge operators, our analysis may create some potential liability in supporting a civil claim of poor stewardship, but at the same time our documentation of the widespread nature of this problem provides a “common practice” defense. Some of the unilateral transfers we have documented, if they were illegitimate (which we do not know), could also create liability for bridge operators, but we believe this potential harm is balanced against the interest of investors in transparency and fairness in operating financial services. Finally, for our proposed real-time analysis approaches, if these were adopted by bridge operators, it could foreclose a broad range of attacks. We admit that this outcome could harm our criminal stakeholders as it might deny them of an important income stream. Knock on effects could include a reduction in revenues for online crypto money laundering services and Lamborghini dealerships, but we believe this is well-justified by the concordant reduction in victimization.
Appendix A Attack Descriptions and Additional Details
In this section we provide a descriptive account of each of the attacks listed in Table 2 in Section 4, giving additional background on the bridge, how the attack transpired, and more details on our results using CrossChecked.
A.1 Ronin Bridge
Background. Ronin bridge operates between Ethereum and the Ronin chain. It was hacked in March 2022. The attacker compromised the bridge’s private keys, allowing them to mint arbitrary amounts of assets. The attacker carried out two transactions, minting around \$624 million USD worth of assets on Ethereum.
Results. CrossChecked analyzed over 3 million bridge transactions and alerted on both attack transactions.
A.2 Poly Network Bridge (2021)
Background. PolyNetwork is a cross-chain bridge that supports asset transfers between multiple blockchains (e.g., BSC, ETH, and Polygon). It was hacked on August 10th, 2021. The attacker exploited a bug in the bridge’s verification code, allowing them to insert their own keys and verify any malicious payload. Overall, the attacker stole around \$600 million USD worth of assets on BSC, ETH, and Polygon.
Results. In total, CrossChecked analyzed over 292k bridge transactions between ETH, BSC, Polygon and Poly Network’s liquidity pool. CrossChecked alerted on all 18 bridge transactions in the attack. In addition, CrossChecked also flagged one withdrawal that seemingly was relayed to the wrong destination chain.
A.3 Binance Token Hub
Background. Binance Token Hub facilitates asset transfers between Binance Beacon Chain and Binance Smart Chain. It was hacked on October 7th, 2022. The attacker exploited a bug in Binance Bridge’s verification code, allowing them to verify carefully crafted malicious payloads. The attacker carried out the attack in two transactions, each minting 1,000,000 BNB on Binance Smart Chain (approximately \$580 million USD in total).
Results. CrossChecked analyzed over 2M bridge transactions between Binance Beacon Chain and Binance Smart Chain. CrossChecked only alerted on the two attack transactions, citing the discrepancy between the amount given out and the amount received by the bridge. We end by noting that the partner deposit transactions returned by the Binance Token Hub’s API for two attack transactions are different than the ones suggested in some blogposts [8]. However, the attack transactions will be flagged regardless of which partner deposit transactions is used.
A.4 Wormhole
Background. Wormhole is a general-purpose cross-chain bridge that currently supports around 35 blockchains. It was hacked on February 2nd, 2022. Specifically, the attacker exploited a bug in Wormhole’s smart contract on Solana that allowed them to verify arbitrary unauthorized payload. The attacker executed one transaction and minted 120,000 wETH (around \$350 million) on Solana.
Results. In total, CrossChecked analyzed over 642k transactions across ten blockchains. It alerted on three transactions, one of which was the attack transaction.
A.5 Nomad Bridge
Background. Nomad bridge support asset transfers across six blockchains. It was hacked on August 1st, 2022. The attacker exploited a bug in the bridge’s verification code, allowing them to verified any payload that had not been verified before. Shortly after the first a few attack transactions, a group of copycats joined the crusade of draining the bridge. In total, the reported loss was around \$190 million.
Results. CrossChecked analyzed over 37k transactions, and alerted on 962 transactions that transferred assets to 561 unique addresses. We found one dataset that reported 561 addresses, which matched exactly with the addresses we identified. We also found one Github repository that mentioned identifying 960 transactions [20]. As mentioned earlier, we identified two more transactions. Upon manually inspection, we confirm that the two transactions were indeed malicious.
A.6 Harmony Bridge
Background. Harmony bridge operates between ETH, BSC and Harmony. It was hacked on June 24, 2022. The attacker compromised two of the signing keys of the bridge, allowing them to mint arbitrary amounts of assets. In total, the attacker minted around \$100 million worth of assets on BSC and ETH in 15 transactions.
Results. CrossChecked analyzed over 336k transactions and alerted on 58 transactions, including all 15 attack transactions.
A.7 HECO Bridge
Background. HECO bridge allows users to transfer assets between Huobi ECO Chain (HECO) and Ethereum. It was hacked on November 11th, 2023. The attacker compromised the bridge’s private keys, allowing them to sign arbitrary transactions. The attacker carried out eight transactions, minting around \$86 million worth of assets on Ethereum.
Results. CrossChecked analyzed over 23k transactions. All eight attack transactions were flagged by CrossChecked.
A.8 Qubit Bridge
Background. Qubit bridge allows users to transfer assets between ETH and BSC. It was hacked on January 27th, 2022. The attacker exploited a bug in the deposit function, which allowed them to trick the bridge into believing that a deposit had been made when it had not. The attacker carried out 16 transactions, stealing around \$80 million worth of assets.
Results. CrossChecked analyzed over 260 transactions and alerted on all 16 transactions.
A.9 Anyswap Bridge
Background. Anyswap bridge supported moving assets across many blockchains at the time of the attack. It was hacked on July 10, 2021. The attacker exploited a bug in the bridge’s verification code, allowing them to verify any malicious payload. The attacker carried out three transactions, minting around \$7.9m worth of assets on Ethereum.
Results. In addition to the previously reported transactions from the July 10, 2021, attack, Anyswap had more than 800 additional transactions that violated the balance invariant.
A.10 Poly Network Bridge (2023)
Background. Shortly after the hack in 2021, PolyNetwork switched to a new set of smart contracts. It was, however, hacked again in August 10th, 2021. The attacker exploited a bug in the bridge’s verification code, allowing them to verify arbitrary payload. Overall, there were 136 reported transactions.
Results. CrossChecked analyzed over 290k transactions and flagged all 136 transactions. CrossChecked also flagged 27 transactions that pointed to the same non-existent deposit transaction hash (0x0101... with 01 repeated 32 times). All of these transactions occurred on Aug 22, 2023 and withdrew funds from the bridges without backing deposits. In total, over \$20 million worth of assets were withdrawn in these transactions.
A.11 Chainswap Bridge
Background. Chainswap bridge support token transfers between five bridges and was hacked on July 10, 2021. The attacker exploited a bug in the bridge’s verification code, allowing them to verify any malicious payload. The attacker stole \$4.4 million worth of assets on Ethereum and BSC using one address. Of particular note, unlike other bridges, Chainswap did not publicly disclose the list of bridge contract addresses. As such, we programmatically identified all potential bridge contract addresses by searching for the specific event signatures that known Chainswap bridge contracts emit.
Results. CrossChecked analyzed over 53k transactions. CrossChecked alerted on 1136 transactions, all of which were executed by the single malicious address. We note that we are unable to find any public information on the list of transactions that were part of the attack, preventing us from verifying individual transactions.
A.12 Meter Bridge
Background. Meter bridge allows users to transfer between Meter’s own chain and a few EVM-based chains. It was hacked on February 5, 2022. The attacker exploited a bug in the bridge’s deposit function, where the attacker tricked the bridge into believing that a deposit had been made. In total, the attacker carried out 5 transactions, stealing around \$4.3 million worth of assets.
Results. CrossChecked analyzed over 14k transactions, alerting on all five attack transactions.
Bridge (Claimed) Deposit Transaction Hash Chain & Token Withdraw Transaction Hash Chain & Token
Wormhole #1 0x542efe5a7f965b927a294bce7f3a30492d0e2ca60da2e2d6cdfaf523a28a2a9c Polygon 0.5 WSOL 5AiesW9pKrZvCCJM8QPWmqx...dkadV1waPVLWfsnCMVQmyYaciLxEo SOL 650 USDC
Wormhole #2 0x8bbb7befd198a5e90297f451fc43a9e90de083289a041c8af94116c785cf496d Polygon 4.3k MATIC 0xbe81895b1c3172fd69b8d4d9bf726edfdc17083876c440f9414ff316999237d7 AVAX 2.3k WMATIC
Anyswap #1 0x01ba4719c80b6fe911b091a7c05124b64eeece964e09c058ef8f9805daca546b N/A 0xf015a6b06a13a08d3499ece17504d14a95d6af3e04ae11f291dca22dbbf6c991 BSC $5*10^{-8}$ USDC
Anyswap #2 0x01ba4719c80b6fe911b091a7c05124b64eeece964e09c058ef8f9805daca546b N/A 0x9e55b7295880dce76aa8af0f3e3f9e36499ae0bdb28088a5924daf29c6132ceb BSC 55k USDC
Anyswap #3 0xe3b0c44298fc1c149afbf4c8996fb92427ae41e4649b934ca495991b7852b855 N/A 0x4f038804d0622d2eab15d21d902a3fdd3bdfb5427bb5fd65b9eb0a41169534be Polygon 100k USDC
Anyswap #4 0x0x0000000000000000000000000000000000000000000000000000000000000000 N/A 0x98aa9e94d4fd0a05c27eb13ac2e699e4426c8dd9d57d04c0fa09cf4951eb2f94 BSC 650k USDC
Anyswap #5 0x0x0000000000000000000000000000000000000000000000000000000000000000 N/A 0xa67ac5dc308142f89409df89dc85e8fab88c575b3adef77fbc8f51858b7bf7cb Polygon 50k USDC
Anyswap #6 0x28b233a4dbda8b4dfae7245b8fff434de95f6dbd101e1a9cb22a95ded1315a16 Fantom 6.4k POPS 0x76bdcfd5ddfa358bf4181556e3b4f1fdd2d648a246bfab91386bdfbd7b76d01f Avalanche 6.4k POPS
0xf0b5568dfd8a4559d30adc9dfc881875210a3b9dfa680d392b33eb1d2cc86cfa Fantom 6.4k POPS 0xc86297f14f32a33232149025d4e8f8e50985d76ac1b7ccaf181501820c0b1cf7 Avalanche 6.4k (any)POPS
Anyswap #7 0x0x0000000000000000000000000000000000000000000000000000000000000000 N/A 0xde790e8dc59d8bae7ebdf89c4b75267a6e0783219b32aebe83e112aac6c299f5 Avalanche 54k USDC
HECO #1 0x6f9d2e82aef87fc649198976974c05d4c540dacca5043ffee619cc33f3ba4cf5 ETH 5m USDT 0x628e878fb723cf0dd838eb956ce78d23b45b130876a625fd4d283e62ac2289f0 HECO 5m USDT
0x6f9d2e82aef87fc649198976974c05d4c540dacca5043ffee619cc33f3ba4cf5 ETH 5m USDT 0x27a1e6a66b6e0fc5fa805f7400dd07397bb92226926868a82afb44154a32128b HECO 5m USDT
Harmony #1 0x559bc92656a6956a5ffe9eea6f14a5d5993520e31a1a08551d5171ad8f658886 BSC 5.3k BUSD 0xdf3bf1a8227ede87d7905c026c3b6a3504cc81399ebd08e1273e1a9dd2c748a9 Harmony 5.3k BUSD
0x559bc92656a6956a5ffe9eea6f14a5d5993520e31a1a08551d5171ad8f658886 BSC 5.3k BUSD 0x304801a2b33585e6867de0c403535588979ce4d2cf41c6922223d3203589c39d ETH $5*10^{-18}$ BUSD
PolyNet. #1 0x0101010101010101010101010101010101010101010101010101010101010101 N/A 0xd6b7f50e974311082eb4b413219f7198cbf897af4e0f2e9202b10c6afe8fa0a2 ETH 491m PLT
Table 3: Other flagged transactions. PolyNet stands for Poly Network. Full transaction hash for Solana: 5AiesW9pKrZvCCJM8QPWmqxsnRoQwHaQmX8NR9a8BFz3pmt2ypW67zgqeRWdkadV1waPVLWfsnCMVQmyYaciLxEo.
Appendix B Additional Transactions Identified
Table 3 provides additional details on specific example bridge transactions that violate the balance invariant and have otherwise not been previously reported. For each example, the table lists the hashes of the paired deposit and withdrawal transactions, the blockchain, and the number of tokens transferred. If a deposit transaction does not exist or is otherwise invalid, we mark its chain and token as N/A.